OpenStudy (anonymous):

If f(x) = 4x - 3, what is f(x)-1?

OpenStudy (defiance):

If f(x) = 4x - 3. We know that we only subtracted 1 more. so, -3 - 1 = -4 f(x)-1 = 4x - 4

OpenStudy (stuck-help):

subtract 1 from each side f(x)-1=4x-4

OpenStudy (anonymous):

thakyou